# Picometer Sensitive Prototype of the Optical Truss Interferometer for LISA

### v1

@@ I Introduction
The OTI is designed to serve as a “plug and play” solution for $1\frac{\mathrm{pm}}{\sqrt{\mathrm{Hz}}}$ dimensional stability testing in the mHz LISA band. We have developed a compact mode matching system, or input stage, that transforms fiber-coupled 1064 nm light into the proper free-space Gaussian mode for the optical cavity within an input stage that is only 4 cm in length. While Lens 1 in Figure 1 serves as a positive focusing lens for mode matching, Lens 2 serves as both the second mode matching lens as well as the cavity input mirror. The front surface of the optic has a concave curvature and a negative focal length, serving as the second mode matching lens, while the backside has a concave 500 mm radius of curvature with a dielectric mirror coating with $R\geq 99.8\%$. The input stage is accompanied with a return stage used only to house the cavity return mirror that is nominally identical to the input mirror such that they form a 70 cm long, symmetric, and impedance-matched cavity. All the internal components of the input and return units are housed in a block of Schott’s ZERODUR®, polished on one side to allow for hydroxide catalysis bonding to a LISA telescope or other test structure [6, 7]. The optics are held in place with an Invar® sleeve, and the mounting is designed to maintain the position of the cavity mirrors with respect to the zerodur housings under thermal variations. Further details on the design, assembly, and alignment of these prototype units can be found in Jersey et al., 2023 [5]. In principle, all that is needed to form a cavity baseline to test the dimensional stability is a suitable test structure and a mounting or bonding method to maintain the alignment between the input and return stages. In this paper, we will discuss the experimental setup and results from our initial performance testing of a first-generation prototype optical truss interferometer.
@@ II.2 Vacuum Chamber Isolation
The free running temperature noise that we measure in the vacuum chamber is typically on par with the LISA requirements above 0.1 mHz. The flattening of the spectrum seen around 1 mHz is a readout limitation due to the analog-to-digital conversion (ADC) used by the multimeter, and the true temperature noise at and above 1 mHz is estimated to be well below the requirement [10]. The noise below 0.1 mHz is above the LISA requirement likely because we allow the temperature to freely run rather than using active stabilization. Even without active temperature control, the cavity spacers and plates we are using, made of zerodur and ULE which have CTE on the order of 10 - 100 ppb/K, are stable enough to still meet the $1\frac{\mathrm{pm}}{\sqrt{\mathrm{Hz}}}$ LISA requirement in this environment. However, it is the OTI input and return stages, as well as the steel mounts holding them to the plates, that may dominate the thermal expansion of the cavities. Thus, it is the effective CTE to which each individual component contributes that will limit the cavity displacement noise at low frequencies where temperature noise dominates [15, 16, 17].
@@ II.3 Cavity Readout Scheme
The foundation of how we measure the displacement noise in the optical truss cavities lies in the Pound-Drever-Hall (PDH) frequency locking method. PDH locking involves creating a feedback loop that maintains the laser frequency on resonance with an optical cavity. By phase modulating the incident field and measuring the field reflected from the cavity, an error signal can be generated, filtered, and fed back into the laser controller to maintain resonance. For the OTI, we employ commercially available fiber-coupled electro-optic phase modulators (EOMs) that are placed in the optical path before coupling into polarization-maintaining optical fibers that feed into the vacuum chamber and into each respective cavity. The reflected field from each cavity is isolated from the incident field via an optical circulator. This is done with a PBS and QWP for the reference cavity, and with commercially available fiber-optical circulators for the OTI cavities. By measuring the reflected power with a photodetector and subsequently demodulating with the RF signal driving the phase modulation, the result is an error signal with a zero-crossing that is linearly proportional to the deviation from resonance [18]. The PDH error signal is filtered through proportional-integral (PI) servos and fed back into the laser controllers, which stabilize the laser frequency to the cavity resonance. While the laser is locked to an optical cavity, the laser frequency is coupled with the optical path length fluctuations by the relation
@@
We employ a spectral analysis approach to estimate the amplitude spectral density (ASD) of the beat frequency time series, which can then be converted to the corresponding cavity length noise via Equation 1 assuming that the variations in the beat frequency are dominated by the OTI laser and not the reference [20, 21]. We employ the LTPDA signal processing toolbox in MATLAB, not only to calculate the ASD of time series data but also to perform many other reliable data analysis tasks in both the frequency and time domain [22]. An example of the estimated ASD of the cavity length noise compared to the benchmark LISA requirement is shown in Figure 6 along with other estimated contributions from the various noise sources that were investigated.
@@ III.1 Temperature Noise
Each noise source can couple into the measured cavity length noise through different mechanisms. An intuitive example is the temperature noise in the vacuum chamber coupling into the cavity length through the effects of thermal expansion. In this case, the coupling factor is the effective CTE of the optical cavity being tested, which can vary depending on the materials used in the cavity assembly, the type of cavity mirrors used, as well as the way the mirrors are bonded or mounted to the cavity spacer [16, 17, 15]. The effective CTE can be measured by taking concurrent measurements of the cavity temperature along with the primary beat note measurement and calculating the coupling factor between the two time series with a least-squares fit, depicted in Figure 7. Both time series are low-pass filtered with a 1 mHz corner frequency, below which the correlation between the temperature and beat note is dominant. Note that in our experiment, we have not actively driven the temperature in the chamber and we use the gradual drift around room temperature over many hours or days to correlate with the beat note measurements. Ideally, we would seek a frequency reference that is external to the vacuum chamber or that is temperature invariant such as a molecular transition, and actively drive the temperature to correlate with the beat note. However, we estimate a room temperature ($\approx$ 20oC in our laboratory) CTE of $-1.25\pm 0.23\times 10^{-6}$ K-1 ($250\pm 50$ nm/K) in OTI 1 and $2.0\pm 0.4\times 10^{-7}$ K-1 ($40\pm 10$ nm/K) in OTI 2, which are both much larger than what we would expect for just the ULE spacers or the zerodur blocks housing the OTI units. Since the effective CTE estimates are well above the CTE of zerodur or ULE-based cavities, it is a safe assumption that our measured values roughly estimate the effective CTE of the OTI cavities and that they are likely limited by the metal components and mounts in the assemblies.
@@ III.3 Residual Amplitude Modulation
Figure 9: Comparison of the RAM noise measured in the cavity reflection at different modulation frequencies. Each term is normalized by the average optical power. The effect of changing the modulation frequency to the free spectral range of the parasitic etalon can be seen in both the RAM noise and the measured cavity noise.
@@
Figure 10: The coherence between the quadrature component of the RAM at the cavity reflection and the cavity noise. The solid curves are from measurements with a modulation frequency of 17.8 MHz and the dashed curves are with a frequency of 28.6 MHz.

### v2

@@ I Introduction
The OTI is designed to serve as a “plug and play” solution for $1\frac{\mathrm{pm}}{\sqrt{\mathrm{Hz}}}$ dimensional stability testing in the mHz LISA band. We have developed a compact mode matching system, or input stage, that transforms fiber-coupled 1064 nm light into the proper free-space Gaussian mode for the optical cavity within an input stage that is only 4 cm in length. While Lens 1 in Figure 1 serves as a positive focusing lens for mode matching, Lens 2 serves as both the second mode matching lens as well as the cavity input mirror. The front surface of the optic has a concave curvature and a negative focal length, serving as the second mode matching lens, while the backside has a concave 500 mm radius of curvature with a dielectric mirror coating with $R\geq 99.8\%$. The input stage is accompanied with a return stage used only to house the cavity return mirror that is nominally identical to the input mirror such that they form a 70 cm long, symmetric, and impedance-matched cavity. All the internal components of the input and return units are housed in a block of Schott’s ZERODUR®, polished on one side to allow for hydroxide catalysis bonding to a LISA telescope or other test structure [6, 7]. The optics are held in place with an Invar® sleeve, and the mounting is designed to maintain the position of the cavity mirrors with respect to the zerodur housings under thermal variations. While using Invar to mount the optics is convenient for ground-based prototypes, it should be noted that the LISA spacecraft must maintain strict magnetic cleanliness and Invar would not be an acceptable choice to include in any flight hardware. Further details on the design, assembly, and alignment of these prototype units can be found in Jersey et al., 2023 [5]. In principle, all that is needed to form a cavity baseline to test the dimensional stability is a suitable test structure and a mounting or bonding method to maintain the alignment between the input and return stages. In this paper, we will discuss the experimental setup and results from our initial performance testing of a first-generation prototype optical truss interferometer.
@@ II.2 Vacuum Chamber Isolation
The free running temperature noise that we measure in the vacuum chamber is typically on par with the LISA requirements above 0.1 mHz. The flattening of the spectrum seen around 1 mHz is a readout limitation due to the analog-to-digital conversion (ADC) used by the multimeter, and the true temperature noise at and above 1 mHz is estimated to be well below the requirement [10]. The noise below 0.1 mHz is above the LISA requirement likely because we allow the temperature to freely run rather than using active stabilization. Even without active temperature control, the cavity spacers and plates we are using, made of zerodur and ULE which have CTE on the order of $10-100\times 10^{-9}$ K-1, are stable enough to still meet the $1\frac{\mathrm{pm}}{\sqrt{\mathrm{Hz}}}$ LISA requirement in this environment. However, it is the OTI input and return stages, as well as the steel mounts holding them to the plates, that may dominate the thermal expansion of the cavities. Thus, it is the effective CTE to which each individual component contributes that will limit the cavity displacement noise at low frequencies where temperature noise dominates [15, 16, 17].
@@ II.3 Cavity Readout Scheme
The foundation of how we measure the displacement noise in the optical truss cavities lies in the Pound-Drever-Hall (PDH) frequency locking method. PDH locking involves creating a feedback loop that maintains the laser frequency on resonance with an optical cavity. By phase modulating the incident field and measuring the field reflected from the cavity, an error signal can be generated, filtered, and fed back into the laser controller to maintain resonance. For the OTI, we employ commercially available nonplanar ring oscillator (NPRO) lasers and fiber-coupled electro-optic phase modulators (EOMs) that are placed in the optical path before coupling into polarization-maintaining optical fibers that feed into the vacuum chamber and into each respective cavity. The reflected field from each cavity is isolated from the incident field via an optical circulator. This is done with a PBS and QWP for the reference cavity, and with commercially available fiber-optical circulators for the OTI cavities. By measuring the reflected power with a photodetector and subsequently demodulating with the RF signal driving the phase modulation, the result is an error signal with a zero-crossing that is linearly proportional to the deviation from resonance [18]. The PDH error signal is filtered through proportional-integral (PI) servos and fed back into the laser controllers, which stabilize the laser frequency to the cavity resonance. While the laser is locked to an optical cavity, the laser frequency is coupled with the optical path length fluctuations by the relation
@@
We employ a spectral analysis approach to estimate the amplitude spectral density (ASD) of the beat frequency time series, which can then be converted to the corresponding cavity length noise via Equation 1 assuming that the variations in the beat frequency are dominated by the OTI laser and not the reference [20, 21]. We employ the LTPDA signal processing toolbox in MATLAB®, not only to calculate the ASD of time series data but also to perform many other reliable data analysis tasks in both the frequency and time domain [22]. An example of the estimated ASD of the cavity length noise compared to the benchmark LISA requirement is shown in Figure 6 along with other estimated contributions from the various noise sources that were investigated.
@@ III.1 Temperature Noise
Each noise source can couple into the measured cavity length noise through different mechanisms. An intuitive example is the temperature noise in the vacuum chamber coupling into the cavity length through the effects of thermal expansion. In this case, the coupling factor is the effective CTE of the optical cavity being tested, which can vary depending on the materials used in the cavity assembly, the type of cavity mirrors used, as well as the way the mirrors are bonded or mounted to the cavity spacer [16, 17, 15]. The effective CTE can be measured by taking concurrent measurements of the cavity temperature along with the primary beat note measurement and calculating the coupling factor between the two time series with a least-squares fit, depicted in Figure 7. Both time series are low-pass filtered with a 1 mHz corner frequency, below which the correlation between the temperature and beat note is dominant. Note that in our experiment, we have not actively driven the temperature in the chamber and we use the gradual drift around room temperature over many hours or days to correlate with the beat note measurements. Ideally, we would seek a frequency reference that is external to the vacuum chamber or that is temperature invariant such as a molecular transition, and actively drive the temperature to correlate with the beat note. However, we estimate a room temperature ($\approx$ 20oC in our laboratory) CTE of $1.25\pm 0.23\times 10^{-6}$ K-1 ($250\pm 50$ nm/K) in OTI 1 and $2.0\pm 0.4\times 10^{-7}$ K-1 ($40\pm 10$ nm/K) in OTI 2, which are both much larger than what we would expect if the dominant contributions were from the ULE spacers and the zerodur blocks housing the OTI units ($\sim 10^{-8}$ K-1). Since the effective CTE estimates are well above the CTE of zerodur or ULE-based cavities, it is a safe assumption that our measured values roughly estimate the effective CTE of the OTI cavities and that they are likely limited by the metal components and mounts in the assemblies.
@@ III.3 Residual Amplitude Modulation
Figure 9: Comparison of the RAM noise measured in the cavity reflection at different modulation frequencies. Each term is normalized by the average optical power. The effect of changing the modulation frequency to the free spectral range of the parasitic etalon can be seen in both the RAM noise and the measured cavity noise. The red curves are from measurements with a modulation frequency of 17.8 MHz and the blue curves are with a frequency of 28.6 MHz.
@@
Figure 10: The coherence between the quadrature component of the RAM at the cavity reflection and the cavity noise. The red curves are from measurements with a modulation frequency of 17.8 MHz and the blue curves are with a frequency of 28.6 MHz.